# Active Light Modulation to Counter Manipulation of Speech Visual Content

### v1

@@ Abstract.
High-profile speech videos are prime targets for falsification, owing to their accessibility and influence. This work proposes Spotlight, a low-overhead and unobtrusive system for protecting live speech videos from visual falsification of speaker identity and lip and facial motion. Unlike predominant falsification detection methods operating in the digital domain, Spotlight creates dynamic physical signatures at the event site and embeds them into all video recordings via imperceptible modulated light. These physical signatures encode semantically-meaningful features unique to the speech event, including the speaker’s identity and facial motion, and are cryptographically-secured to prevent spoofing. The signatures can be extracted from any video downstream and validated against the portrayed speech content to check its integrity. Key elements of Spotlight include (1) a framework for generating extremely compact (i.e., 150-bit), pose-invariant speech video features, based on locality-sensitive hashing; and (2) an optical modulation scheme that embeds ¿200 bps into video while remaining imperceptible both in video and live. Prototype experiments on extensive video datasets show Spotlight achieves AUCs $\geq$ 0.99 and an overall true positive rate of 100% in detecting falsified videos. Further, Spotlight is highly robust across recording conditions, video post-processing techniques, and white-box adversarial attacks on its video feature extraction methodologies.
Work done while at Columbia University
copyright: none
@@ 1. Introduction
In the early days of video technology, high-profile speeches were some of the first events to be shared over the new communication medium (first_video_speeches, ). Influential figures capitalized on its unique persuasive power (Donovan_Scherer_2005, ; presidents_and_tv, ), and ever since, video has been a staple of information exchange. Today, this exchange faces a flood of falsified videos of high-profile speeches spreading disinformation and discord.
@@
This paper focuses on addressing falsification of two salient aspects of a speech event: the speaker’s identity and her lip and face movements, which are directly tied to speech content and speed. These elements are particularly persuasive and semantically-rich (sudar2021, ) and have been targeted in numerous incidents (nancy19, ; hadid23, ; swinney2024, ; biden23_draft, ; biden24_lipsync, ; harris24_slow, ; harris24_slur, ; pelosi_trump_tweet, ). Today, realizing these falsifications is easier than ever. An attacker may use any open-source model or online deepfake tool to generate videos of her victim making fabricated statements (hadid23, ; swinney2024, ; biden23_draft, ; biden24_lipsync, ; harris24_slur, ). Even simple edits achievable on most smartphones, such as changing playback speed or splicing clips, can greatly alter a portrayal (e.g., widely-circulated videos claiming to show Nancy Pelosi and Kamala Harris delivering speeches intoxicated (pelosi_trump_tweet, ; nancy19, ; harris24_slow, )). Once disseminated online, these videos blur the lines between fake and real, posing grave political, financial, and social risks. Current technologies for detecting falsification of speech videos have failed to combat these threats.
@@
Existing technologies and their limitations can be characterized based on the stage of a video’s lifetime they initiate protection (Figure 1). A majority of these methods are passive, aiming to detect falsifications after the video is published by identifying artifacts introduced by editing or deepfake models (spotfornow, ; agarwal2020phoneviseme, ; ciftci2020fakecatcher, ). These visual imperfections, however, are diminishing with advances in generative AI, and passive detection methods are increasingly bypassed (pulsedit, ; carlini2020evading, ).
@@
Figure 1. We combat falsification of live speech videos by initiating protection at the earliest possible stage of a video’s lifetime: the physical scene of the speech.
@@
In response, efforts have increasingly shifted to protecting videos at their digital creation. Such methods either incorporate watermarks into generated content (synthid, ) or tag videos with digitally signed credentials upon capture (truepic, ; liu24, ; liu2022vronicle, ; leica_cam, ), generation, or editing (adobe_cai, ). While promising, these approaches face several practical barriers stemming from their reliance on cooperation from all recording parties and video creators. (1) Methods adding verification information to fake content are unlikely to be adopted by malicious parties. With deepfake and editing technologies increasingly democratized, attackers can easily create videos lacking watermarks or credentials and then claim they are real. (2) Capture-time tools require use of specialized apps (truepic, ) or hardware (liu24, ; liu2022vronicle, ; leica_cam, ), which cannot be guaranteed in the large, unregulated audiences of public speech events. (3) Because digital signatures are bound to a video’s low-level representation (i.e., its pixel values), they must be regenerated when any post-processing techniques are applied. These techniques, such as compression and transcoding, are exceedingly common in today’s video sharing workflows. Digital signature methods thus assume user cooperation at each stage of a video’s lifetime. As we see today, without such uniform compliance, an indistinguishable mix of real and fake unsigned content is produced, sowing public confusion about the underlying events.
@@
This work studies a complementary physical approach to the protection of speech videos, seeking to shift protection agency from recording parties to speakers themselves. We envision speakers deploying a device that creates signatures physically at the speech site, such that they are naturally embedded into all real recordings. Such physical signatures encode semantically-meaningful features that are unique to the event and consistent across recording device and position (e.g., representations of speaker lip motion). Legitimate videos inherently pass validation against the signature, while falsified videos possess diverging features and are thus detected. Further, physical signatures are cryptographically-secured, preventing their forgery. While digital signature methods require all parties to independently sign their videos, here, cryptographic data is generated only once, on the deployed physical signature creation device.
@@
Figure 2. Spotlight protecting a live speech. The low-cost core unit creates signatures encoding the event and embeds them into all videos by projecting imperceptible light with a spatial light modulator (SLM).
@@
Figure 3. Overview of Spotlight’s modules and workflow, including: (1) the speech video signature creation; (2) optical signature embedding during the live speech; (3) integrity verification of a later published video of the speech by the video verification module.
@@
Physical signatures provide several benefits in the context of live speeches, owing to their uniquely early initiation of protection. (1) They inherently protect all videos at the event without demanding recording device cooperation, thus creating a canonical version of the speech reinforced by each filming attendee. (2) Since physical signatures capture higher-level features rather than low-level pixel values, they remain valid after benign edits that preserve the video’s semantic content (e.g., compression).
@@
This paper demonstrates Spotlight, a physical signature platform that disseminates signatures via modulated light at the speech site (Figure 2). A speaker places a low-cost Spotlight core unit, serving as trustworthy third-party witness, at the site. The core unit continually extracts semantically-meaningful and robust visual features specific to the speaker’s identity and face and lip motion (referred to as a digest) and generates a message authentication code (MAC) for the digest using its private key. The digest and its MAC make up the signature. Spotlight then encodes the signature data as optical modulations that remain imperceptible both live and in videos, supporting the platform’s broader adoption. These optical modulations nonetheless manifest in all recordings as decodable pixel-level changes. A published video can be verified at any point in its lifetime by extracting the optical signatures and comparing recovered digests to those computed on the portrayed speech event.
@@
Two main technical challenges arise in realizing Spotlight. (1) The limited frame rates of typical cameras result in a low embedding data capacity (hundreds of bits per second). Thus, digests must be highly compact yet highly descriptive and consistent across camera positions. (2) Optical signature embedding must balance competing objectives of imperceptibility, robustness, and data capacity.
@@
We address these challenges with the following contributions. (1) We propose a framework based on locality-sensitive hashing (LSH) to compress pose-invariant, semantically-meaningful speech video feature vectors to just 150 bits while preserving their performance. This framework supports diverse feature vectors, of arbitrarily high-dimensionality, and is independent of the signature dissemination modality. (2) We design a spatio-temporal light modulation scheme that boosts bandwidth across all RGB cameras while remaining imperceptible and resilient against common video post-processing techniques. To the best of our knowledge, this is the first scheme for embedding invisible information into videos from within the environment, taking an important step towards practical physical signatures. (3) We fabricate a Spotlight prototype and examine its performance on 257 minutes of live speeches captured with our core unit deployed and over 1,300 pairs of real and deepfaked videos, spanning varied falsification granularities. Additionally, we assess its imperceptibility and robustness in diverse recording scenarios. (4) We evaluate Spotlight’s robustness against extensive countermeasures, including two sophisticated white-box adversarial attacks aiming to create falsified content that evades detection.
@@
  • Spotlight’s LSH framework supports over 100-fold reduction in the representation size of generic speech video features while maintaining their verification performance.
  • Spotlight attains Area Under Curves (AUCs) $\geq$ 0.99 and a true positive rate of 100% in detecting falsifications of speaker identity and face and lip motion. In challenging scenarios where as little as 1.35 s of a video is modified, Spotlight achieves an AUC $\geq$ 0.90, a 40% gain over the best passive detector baseline evaluated.
  • Spotlight’s semantically-meaningful digests are robust to varied countermeasures and white-box adversarial attacks.
  • Spotlight supports video recording with any RGB camera, at viewing angles up to 60°$\degree$ and distances up to 3 m, even when videos are captured with no optical zoom.
  • Spotlight achieves error-free signature data extraction and verification of videos recorded in extensive indoor and outdoor environments, as well as after common video post-processing methods such as compression, transcoding, and filter application.
  • User studies and LPIPS scores confirm our optical signatures are imperceptible live and in-video in varied deployment scenarios.
@@ 2.1. Video Falsification
This paper focuses on combating visual falsifications of speaker identity and face and lip motion, which determine delivered content. Such falsifications can be made using traditional techniques such as framerate modification, trimming, and cropping. Increasingly, they are achieved via the following deepfake techniques. (1) Face reenactment uses a source video to drive the facial movements of a target image. For example, an attacker may use frames from a legitimate speech video as their targets to create reenactment deepfakes modifying the speaker’s lip movements to match audio of falsified statements (biden24_lipsync, ). (2) Identity/face swaps replace a victim’s face with that of another identity while maintaining the victim’s facial motion and expression; (3) Complete face synthesis generates fictitious faces (i.e., non-existent identities). For speech video falsification, this form of deepfake has the same end effect as an identity swap deepfake, as the perceived identity of the speaker is changed.
@@ 2.2. Preventing and Detecting Fake Videos
We organize existing techniques for preventing and detecting falsified and fake videos into four categories as below. We discuss mechanisms exclusively targeting AI-generated videos (i.e., deepfakes), as well as broader media authenticity initiatives.
@@
Unfortunately, passive detectors are increasingly evaded and even hijacked to improve the quality of fake videos. For instance, (pulsedit, ) bypasses remote photoplethysmography (rPPG)-based detection methods (ciftci2020fakecatcher, ) by generating faces with realistic rPPG signals. Adversarial attacks have conquered several passive detectors (carlini2020evading, ).
@@
Digital signing and watermarking   These active techniques add information to digital content at its creation to enable immediate or downstream verification. Emerging frameworks append cryptographically-signed provenance metadata to files at recording time (liu24, ; truepic, ; leica_cam, ) or upon editing (liu2022vronicle, ; adobe_cai, ). Digital watermarks directly embed verification signals into real (neekhara2022facesigns, ; alattar2020system, ; qureshi2021detecting, ; wang2021faketagger, ; yang2021faceguard, ; masmoudi2022new, ) or synthetic (synthid, ) media. Overall, these digital methods require the cooperation of all video sources throughout the information ecosystem, since they add information on a per-video level. Unfortunately, adversaries are deincentivized from participating in these frameworks and can find alternatives to create synthetic media lacking such information.
[…]
@@
Liveness detection with light   A related line of work employs active illumination from screens to verify that video chat participants are real (gerstner2022detecting, ; shangrealtime, ; livescreen, ; facerevelio, ). These works follow a challenge-response model, where facial appearance is analyzed with respect to dynamic screen illumination. They are thus constrained to video chat scenarios.

### v2

# Combating Falsification of Speech Videos with Live Optical Signatures (Extended Version)

@@ Abstract.
High-profile speech videos are prime targets for falsification, owing to their accessibility and influence. This work proposes VeriLight, a low-overhead and unobtrusive system for protecting speech videos from visual manipulations of speaker identity and lip and facial motion. Unlike the predominant purely digital falsification detection methods, VeriLight creates dynamic physical signatures at the event site and embeds them into all video recordings via imperceptible modulated light. These physical signatures encode semantically-meaningful features unique to the speech event, including the speaker’s identity and facial motion, and are cryptographically-secured to prevent spoofing. The signatures can be extracted from any video downstream and validated against the portrayed speech content to check its integrity. Key elements of VeriLight include (1) a framework for generating extremely compact (i.e., 150-bit), pose-invariant speech video features, based on locality-sensitive hashing; and (2) an optical modulation scheme that embeds ¿200 bps into video while remaining imperceptible both in video and live. Experiments on extensive video datasets show VeriLight achieves AUCs $\geq$ 0.99 and a true positive rate of 100% in detecting falsified videos. Further, VeriLight is highly robust across recording conditions, video post-processing techniques, and white-box adversarial attacks on its feature extraction methods. A demonstration of VeriLight is available at https://mobilex.cs.columbia.edu/verilight.
Video manipulation; deepfake detection; media authenticity
journalyear: 2025copyright: ccconference: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 ACM SIGSAC Conference on Computer and Communications Security (CCS ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3719027.3765112isbn: 979-8-4007-1525-9/2025/10ccs: Security and privacyccs: Computing methodologies Computer vision
@@ 1. Introduction
In the early days of video technology, high-profile speeches were some of the first events to be shared over the new communication medium (Vlku, 2002). Influential figures capitalized on its unique persuasive power (Donovan and Scherer, 2005), and video has since been a staple of information exchange. Today, this exchange faces a flood of falsified videos of high-profile speeches created to spread disinformation and discord.
@@
This paper focuses on combating falsification of two salient aspects of a speech event: the speaker’s identity, and their lip and face movements, which are directly tied to speech content. These elements are particularly persuasive and semantically-rich (Sundar et al., 2021) and have been targeted in numerous incidents (nan, 2019; Wade, 2023; Xander Elliards, 2024; bid, 2023; McCarthy, 2024; har, 2024b, a). Today, realizing these falsifications is easier than ever. An attacker may use any open-source model or online deepfake tool to generate fictitious videos of their victim. Even simple edits achievable on most smartphones, such as changing playback speed, can greatly alter a portrayal (e.g., widely-circulated videos claiming to show Nancy Pelosi and Kamala Harris delivering speeches intoxicated (nan, 2019; har, 2024b)). Once disseminated online, these videos create serious political, financial, and social risks. Unfortunately, current technologies for detecting falsified speech videos have failed to counter these threats.
@@
Existing technologies can be characterized based on the stage of a video’s lifetime they initiate protection (Figure 1). A majority are passive, operating only after videos are published by searching content for artifacts introduced during falsification (Wang et al., 2020; Ciftci et al., 2020). These visual flaws, however, are diminishing with advances in generative AI, and passive detectors are increasingly evaded (Chen et al., 2022; Carlini and Farid, 2020).
@@
Figure 1. We initiate protection of speech videos at the earliest possible stage of a video’s lifetime: the physical scene of the speech.
@@
In response, efforts have increasingly shifted to protecting videos at their digital creation. Such methods either watermark generated content (syn, 2024) or tag videos with digitally-signed credentials upon capture (tru, 2025; Liu et al., 2024, 2022; Lyon, 2023), generation, or editing (ado, 2024). While promising, these approaches face several barriers stemming from their reliance on cooperation from all recording and editing parties. (1) Methods adding verification information to fake content are unlikely to be adopted by malicious actors. With deepfake and editing technologies increasingly democratized, attackers can easily create videos lacking watermarks or credentials and then claim they are real. (2) Capture-time approaches require the use of specialized apps (tru, 2025) or hardware (Liu et al., 2024, 2022; Lyon, 2023), which cannot be guaranteed in the large audiences of public speech events. (3) Because digital signatures are bound to a video’s low-level representation (i.e., its pixel values), they must be regenerated when any post-processing techniques are applied. These techniques, such as compression and transcoding, are exceedingly common in video sharing workflows. Digital signature methods thus assume user cooperation at each stage of a video’s lifetime. As evident today, without such uniform compliance, an indistinguishable mix of real and fake unsigned content is produced, creating competing narratives of the truth.
@@
This work studies a complementary physical approach to protecting speech videos, seeking to shift protection agency from recording parties to speakers themselves. We envision speakers deploying a device that creates signatures physically at the speech site, such that they are naturally embedded into all real recordings. These physical signatures encode semantically-meaningful features (e.g., representations of speaker lip motion) that are unique to the event and consistent across recording device and position. Legitimate videos inherently pass validation against the signature, while falsified videos possess diverging features and are thus detected. Further, physical signatures are cryptographically-secured, preventing their forgery. While digital signature methods require all parties to independently sign their videos, here, cryptographic data is generated only once, on the deployed physical signature creation device. [^1] We use the term physical signature to convey the use of secure, auxiliary data to verify content. Physical signatures are not, however, created in the same way as digital signatures.
@@
Figure 2. VeriLight protecting a live speech. The low-cost core unit creates signatures encoding the event and embeds them into all videos by projecting imperceptible light with a spatial light modulator (SLM).
@@
Figure 3. Overview of VeriLight’s modules and workflow, including: (1) speech video signature creation; (2) optical signature embedding during the live speech; (3) verification of a later published video of the speech by the video verification module.
@@
Physical signatures provide several benefits in the context of live speeches. (1) They inherently protect all videos at the event without demanding device cooperation, thus creating a canonical version of the speech reinforced by each recording attendee. (2) Since physical signatures capture higher-level features rather than low-level pixel values, they remain valid after benign edits that preserve the video’s semantic content (e.g., compression).
@@
This paper demonstrates VeriLight, a physical signature platform that disseminates signatures via modulated light at the speech site (Figure 2). A speaker places a low-cost VeriLight core unit, serving as a trustworthy third-party witness, at the site. The core unit continually extracts semantically-meaningful and robust visual features specific to the speaker’s identity and face and lip motion (referred to as a video descriptor), and uses its private key to generate a message authentication code (MAC). The descriptor and its MAC make up the signature. VeriLight then encodes the signature data as optical modulations that remain imperceptible both live and in videos, supporting the platform’s broader adoption. These optical modulations nonetheless manifest in all recordings as decodable pixel-level signals. A published video can be verified at any point in its lifetime by extracting the optical signatures and comparing recovered descriptors to those computed on the portrayed speech.
@@
Two main technical challenges arise in realizing VeriLight. (1) The limited frame rate of typical cameras results in a low embedding data capacity (hundreds of bits per second). Thus, descriptors must be highly compact yet highly expressive and consistent across camera positions. (2) Optical signature embedding must balance conflicting goals of imperceptibility, robustness, and data capacity.
@@
We address these challenges with the following contributions. (1) We propose a framework based on locality-sensitive hashing (LSH) to compress pose-invariant, semantically-meaningful speech video feature vectors to just 150 bits while preserving their performance. This framework supports diverse feature vectors and is independent of the signature dissemination modality. (2) We design a spatio-temporal light modulation scheme that boosts bandwidth across all RGB cameras while remaining imperceptible and resilient against common video post-processing techniques. To the best of our knowledge, this is the first scheme for embedding invisible information into videos from within the environment, taking an important step towards practical physical signatures. (3) We fabricate a VeriLight prototype and examine its performance on 257 minutes of speech video captured with our core unit deployed and over 1,300 pairs of real and deepfaked videos. Additionally, we assess its imperceptibility and robustness in diverse recording scenarios. (4) We evaluate VeriLight’s robustness against extensive countermeasures, including two white-box adversarial attacks aiming to generate fake videos that evade detection.
@@
  • VeriLight’s LSH framework supports over 100-fold reduction in the representation size of generic speech video features while maintaining their verification performance.
  • VeriLight attains Areas Under Curves (AUCs) $\geq$ 0.99 and a true positive rate of 100% in detecting falsifications of speaker identity and face and lip motion. In challenging scenarios where as little as 1.35 s of a video is modified, VeriLight achieves an AUC of $\geq$ 0.90, a 40% gain over the best passive detector baseline evaluated.
  • VeriLight’s semantically-meaningful video descriptors are robust to varied countermeasures and white-box adversarial attacks.
[…]
  • VeriLight achieves error-free signature data extraction and verification of videos recorded in extensive indoor and outdoor environments, as well as after common video post-processing methods such as compression, transcoding, and filter application.
  • User studies and perceptual metrics confirm optical signatures are imperceptible live and in video in varied deployment settings.
Artifact availability   Our code and guides on reproducing our hardware are available on Github (ver, 2025a) and Zenodo (ver, 2025b).
@@ 2.1. Video Falsification
This work combats visual falsifications of speaker identity and face and lip motion, which determine delivered content. Such falsifications can be made using traditional techniques like frame rate modification, trimming, or cropping. Increasingly, they are achieved via the following types of deepfakes. (1) Face reenactments use a source video to drive the facial movements of a target image. For example, an attacker may use frames from a real speech video as targets to create reenactment deepfakes wherein the speaker’s lips are reanimated to match fabricated audio (e.g.,  (McCarthy, 2024)). (2) Identity swaps replace a victim’s face with that of another identity while retaining the victim’s facial motion and expression; (3) Complete face synthesis generates fictitious faces (i.e., non-existent identities). In our case, this form of fake has the same end effect as an identity swap deepfake, as it changes the perceived identity of the speaker.
@@ 2.2. Preventing and Detecting Fake Videos
We organize existing techniques for preventing and detecting falsified videos into four categories as below. We discuss mechanisms exclusively targeting AI-generated videos (i.e., deepfakes), as well as broader media authenticity initiatives.
Passive detectors   These techniques analyze videos for evidence of tampering and typically are designed to detect deepfakes. They hone in on high-level physical inconsistencies (Li et al., 2018; Hu et al., 2021; Ciftci et al., 2020; Qi et al., 2020; Agarwal and Farid, 2021; Mittal et al., 2020; Agarwal et al., 2020b; Chugh et al., 2020; Haliassos et al., 2021; Zhou and Lim, 2021) (e.g., unnatural lip movements), biometric incongruities (Agarwal et al., 2023, 2020a, 2019; Cozzolino et al., 2021) (e.g., lack of identity-specific head movements), or pixel-level artifacts learned by models (often convolutional neural networks) trained on deepfakes (Frank et al., 2020; Zhang et al., 2019; Li et al., 2019; Lugstein et al., 2021; Luo et al., 2021; Liu et al., 2021; Yan et al., 2023a; Xu et al., 2023; Wang et al., 2023a, 2020; Afchar et al., 2018; Rossler et al., 2019; Zhao et al., 2021; Chen et al., 2021; Bonettini et al., 2021; Masi et al., 2020; Cao et al., 2022; Nguyen et al., 2019).
@@
Unfortunately, passive detectors are increasingly evaded and even hijacked to improve the quality of fake videos. For instance, (Chen et al., 2022) bypasses remote photoplethysmography (rPPG)-based detection methods (Ciftci et al., 2020) by generating faces with realistic rPPG signals. Adversarial attacks have conquered several passive detectors (Carlini and Farid, 2020).
@@
Digital signing and watermarking   These active techniques add information to digital content at its creation to enable immediate or downstream verification. Emerging frameworks append cryptographically-signed provenance metadata to files at recording time (Liu et al., 2024; tru, 2025; Lyon, 2023) or upon editing (Liu et al., 2022; ado, 2024). Digital watermarks directly embed verification signals into real (Neekhara et al., 2022; Qureshi et al., 2021; Wang et al., 2021) or synthetic (syn, 2024) media. Overall, these digital methods require the cooperation of all video sources throughout the information ecosystem, since they add data on a per-video level. Unfortunately, adversaries are deincentivized from participating in these frameworks and can find alternatives to create synthetic media lacking such data.
Live QR codes   Two prior works display dynamic QR codes by a speaker to disseminate verification information. Both only address falsification of speech audio. Critch (Critch, 2022) proposes to display QR codes that encode a speech’s transcript but does not evaluate the concept. In (Shahid and Roy, 2023), time-frequency features of audio signals are encoded. Because of these features’ lower-level nature, they lack robustness across key recording conditions (e.g., their accuracy is below 90% at distances beyond $2\text{\,}\mathrm{f}\mathrm{t}$ and in the presence of ambient noise) and are not shown to be pose-invariant or adversarially robust. Our work instead focuses on visual falsifications and proposes a generic methodology for compressing pose-invariant, semantically-meaningful features. Our features are robust against diverse ambient conditions and evaluated adversarial attacks.
Independent of the verification features they carry, QR codes impose a strict tradeoff between obtrusiveness and supported camera positions. Prior works show a QR code must be over 20 x 20 cm in size to ensure it is readable by all devices within $3\text{\,}\mathrm{m}$ and 45° (Caria et al., 2019). While using optical zoom can boost recording range, this simply fills more of the view with the QR code. Thus, QR code-based systems force speakers to accept either reduced protection robustness or large flickering patterns at the scene – a critical barrier to adoption.
Light-based methods   A concurrent work (Michael et al., 2025) combats video falsification by modulating scene illumination with secret, pseudo-random signals. By correlating video with these signals, the authors generate images that visualize tampering. These visualizations have the advantage of capturing a broad range of spatial and temporal manipulations, but they must be manually assessed by a knowledgeable analyst. Further, because the encoded signals are not specific to scene content, the method is susceptible to replay attacks, wherein valid signals from real videos may be copied onto fakes. VeriLight prevents such attacks by encoding cryptographically-secure, adversarially robust, and speech-specific features in modulations. It automates video analysis and outputs concrete integrity decisions.
@@
Another related line of work employs active illumination from screens to verify that video chat participants are real (Gerstner and Farid, 2022; Shang and Wu, 2020; Liu et al., 2020; Farrukh et al., 2020). These systems follow a challenge-response model, where facial appearance is analyzed with respect to dynamic screen light. They are thus constrained to video chat scenarios.